Test the divergence theorem for the function v = (xy) X+ (2yz) y+ (3zx) Z. Take as your volume the cube shown in Fig. 1.30, with sides of length 2.


1
Expert's answer
2021-03-17T16:58:12-0400
"V=\int_0^2\int_0^2\int_0^2(y+2z+3x)dxdydz\\V=2\int_0^2\int_0^2(y+2+3x)dxdy
\\V=2^2\int_0^2(3+3x)dx\\V=48\\
V_1=\int_0^2\int_0^2\int_0^2(y)dydzdx=8\\
V_2=\int_0^2\int_0^2\int_0^2(3x)dxdydz=24\\V_3=\int_0^2\int_0^2\int_0^2(2z)dxdydz=16\\8+24+16=48"